# Recurrent neural network analysis of single trajectories switching between anomalous diffusion states

### v1

@@ 1 Introduction
Einstein’s description of diffusion was fundamental in demonstrating that matter is made of atoms by observing their indirect effects on the motion of a tracer [3]. His efforts to learn the physical properties of a system by observing the statistics of a tracer immersed in it were later developed into what became the field of microrheology [8]. Einstein’s assumptions for diffusive processes are very well satisfied for dilute suspensions as the pollen grains observed by Robert Brown [9] and for the motion of colloids in water, free or confined by optical tweezers [10, 11]. However, these assumptions are violated in crowded nonequilibrium environments such as the cell interior, in the motion of microswimmers, and in many active matter and turbulent systems. If any of the assumptions is violated, under general conditions [12, 6, 13, 14], the mean squared displacement grows with a power-law
@@
where $\alpha$ is the anomalous diffusion exponent and $K$ is the generalised diffusion coefficient with dimension $[\mbox{length}^{2}\,\mbox{time}^{-\alpha}]$ [15, 16, 17, 18]. The exponent $\alpha$ discriminates between normal diffusion ($\alpha=1$) and anomalous diffusion ($10<\alpha<$ 1 for sub-diffusion, and $21<\alpha<$ 2 for super-diffusion). Determining the anomalous exponent $\alpha$ and the generalised diffusion coefficient $K$ provides crucial information on the system being studied and sheds light on the physical processes at play, which are not directly observable, as, for example, in studies of chromatin dynamics and genome search processes [19, 20, 21, 22, 23]. It is also of great importance to detect changes in diffusive states and to determine the time at which they occur. Unfortunately, in most systems of interest, it is challenging to experimentally obtain many trajectories under the same conditions [24], which makes it difficult to apply methods relying on ensemble averages such as [25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 13, 35]. This difficulty motivated the development of methods which work on an individual trajectory level [35, 32, 36, 37, 38, 39, 40, 41, 42, 31, 43, 44]), which, however, struggle when only short trajectories are available, a commonplace scenario in non-equilibrium and biophysical experiments. The importance of inferring the anomalous exponent and the generalised diffusion coefficient, combined with difficulty of the task, called for the development of machine-learning approaches (see, e.g.[45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69]), with recent studies developing methods that learn to generate stochastic processes that closely resemble anomalous diffusion ones [70]. In 2020, the Anomalous Diffusion Challenge (AnDi) gathered the community to identify the best methods for analysing anomalous diffusion from single trajectories [71]. Machine-learning approaches proved to be the best-performing ones on single trajectories, a trend confirmed also by the 2024 Anomalous Diffusion Challenge (AnDi2024) [72, 73], which we will discuss in detail below.
@@
Inspired by AnDi2024, in this contribution, we tackle the challenging task of inferring the anomalous diffusion exponent $\alpha$ and generalised diffusion coefficient $K$ from a short single trajectory that displays multiple transitions between different diffusive states, characterised by different $\alpha$ and $K$ [72]. Building on previous work [46, 74, 57] we develop a method based on recurrent neural networks employing Long Short-Term Memory architectures (LSTM) [75, 76]. Following the approach of Ref. [60], our method returns a prediction of $\alpha$ and $K$ as a function of time. We also develop a neural network to detect changes between diffusive states and identify the time at which it occurred, i.e. the changepoint (CP). We also discuss the possibility of inferring a CP from the time-dependent estimate of $\hat{\alpha}(t)$ and $\hat{K}(t)$ using a state-of-the-art changepoint detection algorithm [77] and contrast the performances of this approach with those of a neural network designed to identify the changepoints. We name our method sequandi, benchmark its performances on the AnDi2024 dataset and make it freely available [78].
@@
The present work was motivated by the problems and data of The Anomalous Diffusion Challenge 2024 (AnDi2024) [72]. AnDi2024 is a challenge where multiple research groups competed to showcase the performances of their methods for the inference of the anomalous diffusion exponent $\alpha$ and generalised diffusion coefficient $K$ from single short two-dimensional trajectories (featuring a number of time steps, $T$, between $20$ and $200$) from a common dataset [^1] Part of the challenge featured the analysis of microscopy videos and ensembles of trajectories, but we focus on the single-trajectory analysis., as the one shown in Fig. 1. A notable difference from the first edition of the challenge [71] is that the trajectories can feature multiple transitions between different diffusive states. A segment between two transitions can be as short as $3$ time steps. Establishing when such transitions occur, i.e., identifying the changepoint (CP) times, $t_{i}^{\textrm{CP}}$, is one of the tasks of the challenge.
@@
The anomalous diffusion exponent $\alpha$ varies between 00 and $2$: $20<\alpha<$ 2. The generalised diffusion coefficient $K$ spans a broad range $10^{-12}\leq K\leq 10^{6}$. Except for the immobile state in the quenched trap model, in the other models there are no restrictions within these ranges on the allowed values of $\alpha$ and $K$ between consecutive diffusive states. We report the performances of our method on a benchmark dataset similar to the one of AnDi2024 [79].
@@ 2 Neural network architecture and training procedure
2.1 Architecture.
We build our architecture on the one proposed in Ref. [57], which performed very well on the first AnDi Challenge [71]. As illustrated in Fig. 1, the network features two layers of Long Short-Term Memory (LSTM) cells [76, 75] enabling it to capture long and short-term correlations, while securing stability in the training [80].
Figure 1: Neural Network Architecture of sequandi. 2D trajectory data is normalised and separated into $M$ blocks $b_{\tau}$ of 3 coordinate increments each. The blocks are then fed into two adjacent long short-term memory (LSTM) layers of dimensions 250 and 50. Next, outputs from the second LSTM layer are passed through a dense TimeDistributed layer, which is adapted depending on the task (inference of $\alpha$, $K$ or $t^{\textrm{CP}}$) and returns a sequence of length $M$, ready for post-processing.
An important feature of recurrent neural networks such as LSTMs is that they can receive as input a sequence of arbitrary length and return another sequence of the same length. This allows us to task the network to analyse trajectories of varying length and to return the time-dependent values of $\alpha$ and $K$ along the trajectory, which is suitable to characterise conditions that frequently switch between different behaviours, as suggested in Ref. [60]. The first LSTM layer has dimension 250 and the second one 50. After the second LSTM layer, the sequence goes through a Time-Distributed layer [81]. This connects the sequence returned by the second LSTM to the output layer. Wishing to preserve the sequential output of the LSTM layers, Keras’ TimeDistributed layer applies a fully connected layer for each block in the input sequence, which are processed independently with a shared weight, returning a 2D vector of length matching the input and of a width dependent on the inference task. This layer and the ability to return the time evolution of $\alpha$ and $K$ are the main innovations to Ref. [57]. The shape of the output layer depends on the specific task, as we will discuss below.
@@
We train the networks on the increments between subsequent trajectory coordinates. For a trajectory of length $T$, this means $T-1$ increments per each of the two spatial dimensions $\Delta x_{t}=x_{t+1}-x_{t}$ and $\Delta y_{t}=y_{t+1}-y_{t}$. When working with neural networks, it is customary to normalise the input data (see, e.g., Ref. [46, 74, 57]). However, a normalisation of the increments with respect to their standard deviation or their maxima can obfuscate information about the generalised diffusion coefficient $K$. Therefore, in addition to a typical standardisation, normalising the $\Delta x_{t}$ and $\Delta y_{t}$ in each trajectory together so their combined mean is zero and variance is 1, we include their standard deviation $\sigma$ in the input vector. When training LSTMs, it is often convenient for the training to group the input time series into blocks. The choice of an optimal block size results from several trade-offs. Since the AnDi2024 data features segments with a minimum length of 3, we choose a block to include 3 increments per spatial dimension. The input trajectory is then reshaped into $M$ blocks of dimension 9222Note that $\sigma$ is added on after every coordinate in the block.,
@@
Andi2024 considered a wide range of model experiments. We therefore opted to use a training set which is mixed and contained an equal representation of each model experiment. This also opens the door to apply our network to generic diffusive trajectories, as our approach does not involve a prior screening to first ascertain the type of Challenge-specific diffusive model. We trained our network on simulated trajectories using the andi-datasets Python library provided by the challenge organisers [82]. When training a network to infer a parameter from an interval, there is often a tendency to underestimate the values close to the upper limit of the interval and overestimate the ones close to the lower limit, giving rise to an S-shaped curve when plotting predictions vs the ground truth (see, e.g.[57]). To mitigate this issue, we slightly biased the training dataset to have more cases with $\alpha$ closer to the extremes of the supported (0, 2)02(0,\,2)( 0 , 2 ) range [see Fig. 9(a)]. While the support for $K$ is vast, $[10^{-12},10^{6}]$, we trained for $K<35$. A high representation of $K\leq 1$ was considered as this regime is paradigmatic of confined states and has a higher relative importance in its evaluation with a logarithmic metric, discussed in the next section [see Fig. 9(b)]. Lastly, due to the nature of our CP detection method, we included sets of trajectories with many CP (that is, short segments). In total, we trained on $2.5\times 10^{6}$ simulated trajectories.
We use a batch size of 32, and a cycling of the training data of 20 epochs, with an early stopping monitor to prevent overfitting. As common with LSTM training, we apply 20% recurrent dropout to both LSTM layers to prevent overtraining [83, 57], and 10% of the data is kept for validation. We use the Adam optimizer with learning rate $l_{R}=10^{-3}$ and numerical stability constant $\epsilon=10^{-7}$ [84].
@@ 3.2 Post processing
3.3 Diffusion inference performances
In Fig. 3, we plot an example from the trajectory illustrated in Fig. 1 of the point-wise sequential predictions. The generalised diffusion coefficient is plotted as $\log{[K(t)+1]}$ [^4] This is the argument for the mean squared logarithmic error (MSLE) used to evaluate the performance in inferring $K$ in AnDi2024 and ensures a linear-like loss for small $K$ and a logarithmic one for large $K$.. To assess the performance of our method, we compute the following metrics for the point-wise dynamic predictions $\hat{\alpha}(t)$ and $\hat{K}(t)$ in a given trajectory with $T-1$ increments
@@ 3.3 Diffusion inference performances
Figure 3: Example of $\hat{\alpha}(t)$ and $\hat{K}(t)$ and ground truth. Results from example trajectory in Fig. 1. (a) The orange line is sequandi’s point-wise prediction of the anomalous exponent $\hat{\alpha}(t)$, and the blue piece-wise constant line is the ground truth $\alpha(t)$. (b) The orange line is based on sequandi’s point-wise prediction of the generalised diffusion coefficient $\hat{K}(t)$, and is $\log(\hat{K}(t)+1)$. The blue piece-wise constant line is the ground truth $\log(K(t)+1)$. The prediction tracks the ground truth with an $\textrm{MAE}_{\alpha(t)}=0.134$ and $\textrm{MSLE}_{K(t)}=0.010$.
We measure these metrics for every trajectory in the benchmark dataset, and then take a mean over all the trajectories’ errors. For this dataset, $\textrm{MAE}_{\alpha(t)}=0.191$ and $\textrm{MSLE}_{K(t)}=0.021$  [^5] The AnDi2024 metrics considered a flattened array of all $t$ across all trajectories in the dataset instead of a per-trajectory average. Using this metric we obtain $\textrm{MAE}_{\alpha(t)}=0.176$ and $\textrm{MSLE}_{K(t)}=0.014$.. We discuss these performances in greater detail in Section 5.

### v2

@@ 1 Introduction
Einstein’s description of diffusion was fundamental in demonstrating that matter is made of atoms by observing their indirect effects on the motion of a tracer [3]. His efforts to learn the physical properties of a system by observing the statistics of a tracer immersed in it were later developed into what became the field of microrheology [8]. Einstein’s assumptions for diffusive processes are very well satisfied for dilute suspensions, as the pollen grains observed by Robert Brown [9] and for the motion of colloids in water, free or confined by optical tweezers [10, 11]. However, these assumptions are violated in crowded nonequilibrium environments such as the cell interior, in the motion of microswimmers, and in many active matter and turbulent systems. If any of the assumptions is violated, under general conditions [12, 6, 13, 14], the mean squared displacement grows with a power-law
@@
where $\alpha$ is the anomalous diffusion exponent and $K$ is the generalised diffusion coefficient with dimension $[\mbox{length}^{2}\,\mbox{time}^{-\alpha}]$ [15, 16, 17, 18]. The exponent $\alpha$ discriminates between normal diffusion ($\alpha=1$) and anomalous diffusion ($0<\alpha<1$ for sub-diffusion, and $1<\alpha<2$ for super-diffusion).
@@
Determining the anomalous exponent $\alpha$ and the generalised diffusion coefficient $K$ provides crucial information on the system being studied and sheds light on the physical processes at play, which are not directly observable, as, for example, in studies of chromatin dynamics and genome search processes [19, 20, 21, 22, 23]. It is also of great importance to detect changes in diffusive states and to determine the time at which they occur. Unfortunately, in most systems of interest, it is challenging to experimentally obtain many trajectories under the same conditions [24], which makes it difficult to apply methods relying on ensemble averages such as [25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 13, 35]. This difficulty motivated the development of methods which work on an individual trajectory level [35, 32, 36, 37, 38, 39, 40, 41, 42, 31, 43, 44]), which, however, struggle when only short trajectories are available, a commonplace scenario in non-equilibrium and biophysical experiments. The importance of inferring the anomalous exponent and the generalised diffusion coefficient, combined with the difficulty of the task, called for the development of machine-learning approaches as reviewed, e.g., in [45, 46]. Methods ranging from random forests [47, 48, 49, 50], gradient boosting [49, 50, 51], extreme learning machines [52], (unsupervised) anomaly detection [53], feeding statistical features in a shallow network [54], to deep-learning approaches [55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 71, 72, 73, 74, 75, 76, 77, 78] were developed. These deep-learning methods leveraged different architectures, such as Long Short-Term Memory (LSTM) [55, 56, 57, 58, 59, 60, 61], Bayesian deep learning based on LSTM [60, 61], convolutional neural networks (CNN) [63, 64, 65, 77, 66, 67, 68, 69], combinations of CNN and LSTM architectures [62], deep feed-forward networks [71], convolutional transformers [74, 75, 79, 72], physics-informed transformers [78] and graph neural networks [76, 73]. Recent studies have also developed methods that learn to generate stochastic processes that closely resemble anomalous diffusion ones [80].
In 2020, the Anomalous Diffusion Challenge (AnDi) gathered the community to identify the best methods for analysing anomalous diffusion from single trajectories [81]. Machine-learning approaches proved to be the best-performing ones on single trajectories. Among these, most methods were based on LSTM architectures [82, 83] inspired by [55]. The success of machine-learning methods was also confirmed by the 2024 Anomalous Diffusion Challenge (AnDi2024) [84, 85, 86], which we will discuss in detail below.
Some of these methods managed to tackle scenarios in which the anomalous diffusion exponent displayed a switch between two values [55, 57, 81]. However, the case where trajectories display multiple switches between different anomalous diffusion exponent $\alpha$ and/or generalised diffusion coefficients $K$ has so far proven prohibitively difficult. To illustrate this difficulty, it is instructive to see how the method proposed in Ref. [55] to deal with trajectories displaying a single change in $\alpha$ struggles with an example trajectory featuring two switches where both $\alpha$ and $K$ vary, as shown in Fig. 3. Short trajectories of this kind were the focus of Andi2024 [84]. We participated in this challenge and we report here on how we tackle the challenging task of inferring the anomalous diffusion exponent $\alpha$ and generalised diffusion coefficient $K$ from a short single trajectory that displays multiple transitions between different diffusive states, characterised by different $\alpha$ and $K$.
@@
Building on previous work [55, 87, 57] we develop a method based on recurrent neural networks employing (LSTM) [82, 83]. Following the approach of Ref. [77], our method returns a prediction of $\alpha$ and $K$ as a function of time. We also develop a neural network to detect changes between diffusive states and identify the time at which it occurred, i.e. the changepoint (CP). We also discuss the possibility of inferring a CP from the time-dependent estimate of $\hat{\alpha}(t)$ and $\hat{K}(t)$ using a state-of-the-art changepoint detection algorithm [88] and contrast the performances of this approach with those of a neural network designed to identify the changepoints. We name our method sequandi, benchmark its performances on the AnDi2024 dataset and make it freely available [89].
@@
The present work was motivated by the problems and data of The Anomalous Diffusion Challenge 2024 (AnDi2024) [84]. AnDi2024 is a challenge where multiple research groups competed to showcase the performances of their methods for the inference of the anomalous diffusion exponent $\alpha$ and generalised diffusion coefficient $K$ from single short two-dimensional trajectories (featuring a number of time steps, $T$, between 202020 and 200200200) from a common dataset [^1] Part of the challenge featured the analysis of microscopy videos and ensembles of trajectories, but we focus on the single-trajectory analysis., as the one shown in Fig. 1. A notable difference from the first edition of the challenge [81] is that the trajectories can feature multiple transitions between different diffusive states. A segment between two transitions can be as short as 333 time steps. Establishing when such transitions occur, i.e., identifying the changepoint (CP) times, $t_{i}^{\textrm{CP}}$, is one of the tasks of the challenge.
@@
The anomalous diffusion exponent $\alpha$ varies between 0 and 222: $0<\alpha<2$. The generalised diffusion coefficient $K$ spans a broad range $10^{-12}\leq K\leq 10^{6}$. Except for the immobile state in the quenched trap model, in the other models there are no restrictions within these ranges on the allowed values of $\alpha$ and $K$ between consecutive diffusive states.
We report the performances of our method on a benchmark dataset similar to the one of AnDi2024 [90]. We choose to focus on data from AnDi2024 as it provides a comprehensive and well-curated benchmark with several advantages. It includes biologically relevant models capturing realistic diffusion behaviours (e.g., confinement, dimerisation). The dataset is inspired by experimental observations featuring a diverse range of diffusion regimes and transitions, notably including changes in both the diffusion coefficient $K$ and anomalous exponent $\alpha$. Additionally, the benchmark dataset provides the ground truth, which, in combination with the fact that multiple research groups have already adopted it, enables objective and quantitative evaluation of the performance.
@@ 2 Neural network architecture and training procedure
When training LSTMs, it is often convenient for the training to group the input time series into blocks. The choice of an optimal block size results from several trade-offs, such as the minimal time resolution required, and the computational and memory cost of the training. Since the AnDi2024 data features segments with a minimum length of 3, we choose a block to include 3 increments per spatial dimension. This means that we slice the input trajectories into successive short segments containing three increments each. We infer $K$, and $\alpha$ for each of these small segments and classify them based on whether they contain a changepoint or not. We train the networks on the increments between subsequent trajectory coordinates. For a trajectory of length $T$, this means $T-1$ increments per each of the two spatial dimensions $\Delta x_{t}=x_{t+1}-x_{t}$ and $\Delta y_{t}=y_{t+1}-y_{t}$. When working with neural networks, it is customary to normalise the input data (see, e.g., Ref. [55, 87, 57]). However, a normalisation of the increments with respect to their standard deviation or their maxima can obfuscate information about the generalised diffusion coefficient $K$. Therefore, in addition to a typical standardisation, normalising the $\Delta x_{t}$ and $\Delta y_{t}$ in each trajectory together so their combined mean is zero and variance is 1, we include their standard deviation $\sigma$ in the input vector.
The input trajectory is then reshaped into $M$ blocks of dimension 9222Note that $\sigma$ is added on after every coordinate in the block.,
@@
2.2 Architecture.
As discussed in the introduction, LSTMs [83, 82], with their ability to capture long and short-term correlations, while securing stability in the training [91], have proven very successful in the study of anomalous diffusion [55, 56, 57, 60, 61, 58, 59]. We base our model on the architecture proposed in Ref. [57], which performed very well on the first AnDi Challenge [81]. As in Ref. [57], we choose the first LSTM to have dimension 250 and the second one 50, and adjust the relevant hyperparameters. In B, we show how this choice, motivated by the existing literature, performs equally well to networks with hyperparameters fine-tuned for this specific task. Our method, sequandi, features a key difference compared to [57]: leveraging the LSTM properties, it returns as output a time series of the same length as the input (in terms of the short segments the trajectories are sliced into).
Figure 1: Neural Network Architecture of sequandi. 2D trajectory data is normalised and separated into $M$ blocks $b_{\tau}$ of 3 coordinate increments each. The blocks are then fed into two adjacent LSTM layers of dimensions 250 and 50. Next, outputs from the second LSTM layer are passed through a dense TimeDistributed layer, which is adapted depending on the task (inference of $\alpha$, $K$ or $t^{\textrm{CP}}$) and returns a sequence of length $M$, ready for post-processing.
This allows us to task the network to analyse trajectories of varying length and to return the time-dependent values of $\alpha$ and $K$ along the trajectory, as in Ref. [77]. Crucially, this key innovation makes it possible to characterise trajectories that frequently switch between different behaviours both in terms of $\alpha$ and $K$. This would be prohibitively difficult for networks developed to study scenarios that display a single change in diffusive behaviours, such as the ones in Refs. [55, 57] (see Fig. 3). In addition to inferring the time-dependent behaviour of $\alpha$ and $K$, we are interested in detecting the presence of a changepoint (CP) in each of the short segments the trajectories are sliced into, in a combined regression-classification procedure. Since the input trajectories have different lengths, it is not straightforward to design an architecture of variable length to classify whether any short segment contains a CP. To this aim, we apply a fully connected layer to each of the sequential outputs returned by the second LSTM layer. This layer of variable length is known as a Time-Distributed layer and is readily available in Keras [92]. Crucially, this layer maintains the sequential nature of the LSTM layers, so that each block in the sequence is not analysed in isolation but keeps track of the correlations present in the trajectory. Alternatively, one may have used a masking procedure, which would require specifying a maximum possible trajectory length. Finally, an output layer is attached at the end, with a specific shape depending on the task, as we will discuss below.
@@
Andi2024 considered a wide range of model experiments. We therefore opted to use a training set which is mixed and contains an equal representation of each model experiment. This also opens the door to apply our network to generic diffusive trajectories, as our approach does not involve a prior screening to first ascertain the type of Challenge-specific diffusive model. We trained our network on simulated trajectories using the andi-datasets Python library provided by the challenge organisers [93]. When training a network to infer a parameter from an interval, there is often a tendency to underestimate the values close to the upper limit of the interval and overestimate the ones close to the lower limit, giving rise to an S-shaped curve when plotting predictions vs the ground truth (see, e.g.[57]). To mitigate this issue, we slightly biased the training dataset to have more cases with $\alpha$ closer to the extremes of the supported (0, 2)(0,\,2)( 0 , 2 ) range [see Fig. 9(a)]. While the support for $K$ is vast, $[10^{-12},10^{6}]$, we trained for $K<35$. A high representation of $K\leq 1$ was considered as this regime is paradigmatic of confined states and has a higher relative importance in its evaluation with a logarithmic metric, discussed in the next section [see Fig. 9(b)]. Lastly, due to the nature of our CP detection method, we included sets of trajectories with many CP (that is, short segments). In total, we trained on $2.5\times 10^{6}$ simulated trajectories.
We use a batch size of 32, and a cycling of the training data of 20 epochs, with an early stopping monitor to prevent overfitting. As common with LSTM training, we apply 20% recurrent dropout to both LSTM layers to prevent overtraining [94, 57], and 10% of the data is kept for validation. We use the Adam optimiser with learning rate $l_{R}=10^{-3}$ and numerical stability constant $\epsilon=10^{-7}$ [95]. See Fig. 10 for the training and validation loss curves.
@@ 3.2 Post processing
3.3 Diffusion inference performance
We analyse the performance of sequandi on two separate large datasets containing trajectories that were not seen by the networks during training: the benchmark dataset and the test dataset. The benchmark dataset was provided by the organisers of Andi2024, it is similar to the one used for the challenge [90], and we discuss it in Section 5.4. The test dataset is statistically similar to the dataset we used to train sequandi, and we discuss it in  A. It is mostly useful to indicate that the training did not lead to overfitting.
To assess the performance of our method, we compute the following metrics for the point-wise dynamic predictions $\hat{\alpha}(t)$ and $\hat{K}(t)$ in a given trajectory with $T-1$ increments
@@ 3.3 Diffusion inference performance
We measure these metrics for every trajectory in the benchmark dataset, and then take a mean over all the trajectories’ errors. For this dataset, $\textrm{MAE}_{\alpha(t)}=0.191$ and $\textrm{MSLE}_{K(t)}=0.021$  [^4] The AnDi2024 metrics considered a flattened array of all $t$ across all trajectories in the dataset instead of a per-trajectory average. Using this metric we obtain $\textrm{MAE}_{\alpha(t)}=0.176$ and $\textrm{MSLE}_{K(t)}=0.014$.. We discuss these performances in greater detail in Section 5. For the test dataset we find $\textrm{MAE}_{\alpha(t)}=0.191$ and $\textrm{MSLE}_{K(t)}=0.046$.
To illustrate our findings, it is also instructive to visualise the predictions made by sequandi on example trajectories. We start by considering a trajectory from the test dataset (shown in Fig. 1), presenting the challenging inference of two switches, in rapid succession, between anomalous diffusion states. In Fig. 3, we plot the point-wise sequential predictions. The generalised diffusion coefficient is plotted as $\log{[K(t)+1]}$ [^5] This is the argument for the mean squared logarithmic error (MSLE) used to evaluate the performance in inferring $K$ in AnDi2024 and ensures a linear-like loss for small $K$ and a logarithmic one for large $K$.. To visualise the improvement provided by sequandi, we compare it to the standard approach of computing a Time-Averaged Mean-Squared Displacement (TA-MSD) on a sliding window (brown line). For $\alpha$ we can also compare it to an adaptation of previous state-of-the-art time-dependent inference of $\alpha$ [55], which gives the light-grey curve of Fig. 3 (see E.1 for details). In E.2, we visualise the performance on more example trajectories.
Figure 3: Example of $\hat{\alpha}(t)$ and $\hat{K}(t)$ and ground truth. Results from example trajectory in Fig. 1. (a) The orange line is sequandi’s point-wise prediction of the anomalous exponent $\hat{\alpha}(t)$, and the blue piece-wise constant line is the ground truth $\alpha(t)$. The grey line is the prediction of the network developed in Ref. [55] and the brown line is from the TA-MSD. Since these methods require longer trajectories to work, the original trajectory was repeated (see E.1 for details). (b) The orange line is based on sequandi’s point-wise prediction of the generalised diffusion coefficient $\hat{K}(t)$, and is $\log(\hat{K}(t)+1)$. The blue piece-wise constant line is the ground truth $\log(K(t)+1)$, and the brown line is obtained via a TA-MSD estimate. sequandi’s prediction tracks the ground truth with an $\textrm{MAE}_{\alpha(t)}=0.134$ and $\textrm{MSLE}_{K(t)}=0.016$. In this window of 200 frames, the error from the network of Ref. [55] is $\textrm{MAE}_{\alpha(t)}=0.268$, while for the TA-MSD they are $\textrm{MAE}_{\alpha(t)}=0.295$ and $\textrm{MSLE}_{K(t)}=0.194$